A 3.72-kg iron box is suspended from a spring with a spring constant of k =...

  1. A 3.72-kg iron box is suspended from a spring with a spring constant of k = 18 N/m. The mass is pulled 0.35 m downward from its equilibrium position and allowed to oscillate.

    1. What will the angular velocity be for this iron box?

    2. With what period will the iron box oscillate?
Answer #1

Angular velocity = sqroot( k/m)

Or angular velocity = sqroot(18/3.72) = 2.1997 rad / sec

Now omega = 2×pi/T

So time period T = 2×3.14/2.1997 = 2.855 sec